For this question my professor was talking fast but the things I caught were 1) working capital 2) work ratio 3) quick ratio that I need to find



Consolidated Balance Sheets Dollars in millions except per share amounts December 31 2016 2017 Assets Current Assets s50,498 16,522 1,369 s 5,788 16,794 1,555 14,232 Cash and cash equivalents Accounts receivable net of allowances for doubtful accounts of $663 and $661 Prepaid expenses 10,757 79,146 125,222 105,449 96,136 10,676 7,464 1,560 18,444 Other current assets 38,369 Total current assets 124,899 105,207 94,176 14,243 8,441 1,674 16,812 Property, Plant and Equipment- Net Goodwil Licenses Customer Lists and Relationships-Net Other Intangible Assets - Net Investments in Equity Affiliates Other Assets Total Assets 444,097 $403,821 Liabilities and Stockholders Equity Current Liabilities Debt maturing within one year Accounts payable and accrued liabilities Advanced billings and customer deposits Accrued taxes Dividends payable Total current liabilities 38,374 34,470 4,213 1,262 3,070 81,389 125,972 31,138 4,519 2,079 3,008 50,576 113,681 Long-Term Debt Deferred Credits and Other Noncurrent Liabilities Deferred income taxes Postemployment benefit obligation Other noncurrent liabilities 43,207 31,775 19,747 94,729 60,128 3.578 21,748 115,454 Total deferred credits and other noncurrent liabilities Stockholders Equity Common stock ($1 par value, 14,000,000,000 authorized at December 31, 2017 and 2016: issued 6,495,231,088 at December 31, 2017 and 2016) Additional paid-in capital Retained earnings Treasury stock (355,806,544 at December 31, 2017 6,495 89,563 50,500 6,495 89,604 34,734 (12,659) 4,961 975 and 356,237,141 at December 31, 2016, at cost) Accumulated other comprehensive income Noncontrolling interest Total stockholders equity Total Liabilities and Stockholders Equity (12,714) 7,017 1,146 142,007T $444,097 124,110 $403,821 notes are an integral part of the consolidated financial statements
Answer #1

1) Working capital = current assets - current liabilities

  1. For 2017 : 79146- 81389= - 2243
  2. for 2016: 38369-50576= -12,207

2) There is no such thing as work ratio. However there is a working capital ratio= current assets /current liabilities. This is the same as current ratio.

  1. For 2017 : 79146/81389= 0.972
  2. for 2016: 38369/50576= 0.759

3) Quick ratio = (cash + marketable securities + accounts receivable) / current liabilities

OR (current liabilities- inventory)/ current liabilities.

OR (cash + cash equivalents +short term investments+ account receivables)/ current liabilities

the third formula has been used to calculate the quick ratio

  1. For 2017: (50498+ 16522)/ 81389= 0.823
  2. for 2016: ( 5788+ 16794) / 50576=0.447